# Online Variational Sequential Monte Carlo

### v1

@@ 1. Introduction
Being the most classical structured probabilistic generative model for serial data, state-space models (SSM), also known as general state-space hidden Markov models, are fundamental and ubiquitous in AI and statistical machine learning (Cappé et al.,, 2005; Bishop,, 2016). In SSM, any form of parameter learning or state inference typically involves the computation of complex joint posterior distributions of latent-state variables—the so-called joint-smoothing distributions—given records of observations, which is a delicate problem whose analytical solution is intractable outside the limited cases of linear Gaussian models or models with finite state space. In a recent line of research (Le et al.,, 2018; Maddison et al.,, 2017; Naesseth et al.,, 2018), this problem is addressed by combining variational inference (Kingma and Welling,, 2014) and sequential Monte Carlo (SMC) methods (Gordon et al.,, 1993; Doucet et al.,, 2001; Chopin and Papaspiliopoulos,, 2020) in order to design flexible families of variational joint-state distributions in the form of particle-trajectory laws. By optimizing the Kullback-Leibler divergence from the law of the particle trajectories to the joint-smoothing distribution, this variational SMC (VSMC) approach is killing two birds with one stone by learning not only the unknown model parameters but also, in parallel, an optimal particle proposal kernel, the latter being a problem that has received a lot of attention in the literature (Doucet et al.,, 2000; Cornebise et al.,, 2008; Gu et al.,, 2015). The procedure can be viewed as a non-trivial extension of the importance-weighted auto-encoder (IWAE) proposed by Burda et al., (2016), where standard self-normalized importance sampling has been replaced by sequential importance sampling with systematic resampling in order to obtain a tighter evidence lower bound (ELBO). The objective of VSMC is a similar ELBO whose gradient is approximated by the expectation, under the law of the random numbers generated by the SMC algorithm, of the gradient of the logarithm of the standard particle-based likelihood estimator, the latter being unbiased (Del Moral,, 2004; Chopin and Papaspiliopoulos,, 2020), is obtained as a by-product of the particle approximation of the marginal-state posterior—or filter distribution—flow; thus, whereas traditional particle-based inference in SSM typically relies on particle approximation of the joint-smoothing distributions (with aim of approximating, e.g., the intermediate quantity of the expectation–maximization, EM, algorithm or the score function directly via the Fisher identity), which is cumbersome due to the so-called particle-path degeneracy phenomenon (Kitagawa,, 1996; Cappé et al.,, 2005, Section 8.3), VSMC allows, using the reparameterization trick, the model parameters as well as an optimal particle proposal kernel to be simultaneously learned by processing repeatedly the given data batch using a standard particle filter.
@@
In its basic form, VSMC is an offline inference technique, in the sense that it requires the full data batch to be processed between every update of the model and variational parameters. Still, in a wide range of AI and machine-learning contexts, observed data become available sequentially through a data stream, requiring learning to be performed in an online fashion. The increasing interest in online machine-learning technology also stems from the need of processing data batches of ever-increasing sizes. Thus, in this work we propose an online, stochastic approximation-based version of VSMC, online variational SMC (OVSMC), which can be used for simultaneous online model learning and proposal adaptation. On the contrary to traditional approaches to particle-based online parameter learning in SSM such as particle-based recursive maximum-likelihood (RML, Le Gland and Mevel,, 1997; Del Moral et al.,, 2015) or online EM (Mongillo and Denève,, 2008; Cappé,, 2011), which rely on particle-based online approximation of the tangent filter (filter derivative) and the EM-intermediate quantity, respectively, our approach does not involve any particle smoothing, which, in order to avoid the problem of collapsing particle ancestral genealogies, typically calls for backward-sampling techniques which can be computationally very costly. In addition, the method allows for effective online adaptation of the particle proposal kernel in a way that differs from typical approaches in the literature; indeed, whereas traditional approaches aim to optimize the proposal time step by time step on the basis of local criteria (see, e.g., Cornebise et al.,, 2008, 2014; Zhao et al.,, 2022), our approach is based on a global KL-based criterion (described in detail in Section 3) which allows the particle cloud to be effectively guided toward state-space regions of high local likelihood, without running the risk of over-adapting locally the proposal to current (or temporarily neighboring) observations.
@@
Although our proposed method has similarities with the streaming variational Monte Carlo methodology proposed by Zhao et al., (2022), it is essentially different from the same. An important difference is that the aforementioned work focuses on local optimization of model and variational parameters, such that these are assumed to vary with time and are therefore optimized time step by time step, while OVSMC estimate global, static parameters using stochastic approximation. In addition, in order to show that our approach is statistically well founded, we provide rigorous theoretical results describing its convergence (Theorem 4.7). Under certain strong mixing assumptions, the time-normalized gradient guiding the learning of batch VSMC can, using Birkhoff’s ergodic theorem, be shown to converge as the observation batch size increases towards infinity to a deterministic function depending on the parameter as well as the particle sample size; we show that as the number of observations tends to infinity, OVSMC is solving the same problem as this ideal, ‘asymptotic’ VSMC, in the sense that the mean field targeted by OVSMC coincides with the time-normalized asymptotic gradient of VSMC.
@@
Finally, we illustrate OVSMC numerically on a number of classical SSM, for which the method exhibits fast parameter learning and efficient adaptation of the particle proposal kernel. In the same numerical study, we also show that OVSMC is a strong challenger of VSMC on batch problems.
@@ 2. Background
(1) $p_{\theta}(x_{0:t},y_{0:t})=m_{0}(x_{0})g_{\theta}(y_{0}\mid x_{0})\prod_{s=0}^{t-1}m_{\theta}(x_{s+1}\mid x_{s})g_{\theta}(y_{s+1}\mid x_{s+1}).$
@@ 2.1. Sequential Monte Carlo Methods
See Del Moral, (2004) for a comprehensive treatment of the theory of SMC. The SMC procedure consists of two core operations performed alternately: a selection step, which resamples the particles with replacement according to their weights, and a mutation step, which propagates randomly the selected particles to new locations. After importance sampling-based initialization of $(\xi_{0}^{i},\omega_{0}^{i})_{i=1}^{N}$, the random sample is updated according to Algorithm 1, in which mutation (Line 3) is executed using some generic proposal Markov transition kernel, possibly depending on $y_{t+1}$, with density $r_{\lambda}(x_{t+1}\mid x_{t},y_{t+1})$ parameterized by some vector $\lambda$ belonging to some parameter space $\mathsf{\Lambda}$. Line 2 corresponds to the selection step, where indices $(I_{t+1}^{i})_{i=1}^{N}$ guiding the resampling are drawn from the categorical distribution on $\{1,\dots,N\}$ induced by the particle weights.
@@
Determining a good proposal distribution $r_{\lambda}$ is crucial for the performance of the particle filter (see, e.g., Cornebise et al.,, 2008, 2014; Gu et al.,, 2015; Zhao et al.,, 2022), and the particles should be guided towards state-space regions of non-vanishing likelihood in order to avoid computational waste. For instance, if the latent process is diffuse while the observations are highly informative, letting naively, as in the so-called bootstrap particle filter (Gordon et al.,, 1993), $r_{\lambda}\equiv m_{\theta}$ may result in many particles being assigned a negligible weight and thus significant sample depletion. A more appealing, data-driven option is to use the locally optimal proposal satisfying $r_{\lambda}(x_{t+1}\mid x_{t},y_{t+1})\propto m_{\theta}(x_{t+1}\mid x_{t})g_{\theta}(y_{t+1}\mid x_{t+1})$; however, this proposal is available in a closed form only in a few cases (Doucet et al.,, 2000; Cappé et al.,, 2005, Section 7.2.2.2). In Section 3 we will present a technique that allows to learn simultaneously, in an online fashion, both unknown model parameters and an efficient proposal. The proposed method is based on variational inference, which is briefly reviewed in the next section.
@@ 2.2. Variational Inference
(3) $\mathcal{L}^{\texttt{VI}}(\lambda,\theta)\vcentcolon=\mathbb{E}_{q_{\lambda}}\left[\log\left(\frac{p_{\theta}(X_{0:T},y_{0:T})}{q_{\lambda}(X_{0:T}\mid y_{0:T})}\right)\right]\leq\log\mathbb{E}_{q_{\lambda}}\left[\frac{p_{\theta}(X_{0:T},y_{0:T})}{q_{\lambda}(X_{0:T}\mid y_{0:T})}\right]=\log p_{\theta}(y_{0:T})$
@@
The importance weighted autoencoder (IWAE, Burda et al.,, 2016) extends this idea further by optimizing a similar but improved ELBO, where the expectation is taken with respect to $N$ independent and $q_{\lambda}$-distributed random variables:
(4) $\mathcal{L}^{\tiny{\texttt{IWAE}}}(\lambda,\theta)\vcentcolon=\mathbb{E}_{q_{\lambda}^{\varotimes N}}\left[\log\left(\frac{1}{N}\sum_{i=1}^{N}\frac{p_{\theta}(X_{0:T}^{i},y_{0:T})}{q_{\lambda}(X_{0:T}^{i}\mid y_{0:T})}\right)\right]\leq\log\mathbb{E}_{q_{\lambda}^{\varotimes N}}\left[\frac{1}{N}\sum_{i=1}^{N}\frac{p_{\theta}(X_{0:T}^{i},y_{0:T})}{q_{\lambda}(X_{0:T}^{i}\mid y_{0:T})}\right]\\ =\log\mathbb{E}_{q_{\lambda}}\left[\frac{p_{\theta}(X_{0:T},y_{0:T})}{q_{\lambda}(X_{0:T}\mid y_{0:T})}\right]=\log p_{\theta}(y_{0:T}).$
@@
$\mathcal{L}^{\texttt{SMC}}(\lambda,\theta)\vcentcolon=\mathbb{E}_{Q^{N}_{\lambda,\theta}}\left[\log\left(\prod_{t=0}^{T}\frac{1}{N}\Omega_{t}\right)\right],~$
@@
(5) $\omega_{t+1}^{i}(\lambda,\theta)\vcentcolon=m_{\theta}(f_{\lambda}(\xi_{t}^{I_{t+1}^{i}},y_{t+1},\varepsilon_{t+1}^{i})\mid\xi_{t}^{I_{t+1}^{i}})\frac{g_{\theta}(y_{t+1}\mid f_{\lambda}(\xi_{t}^{I_{t+1}^{i}},y_{t+1},\varepsilon_{t+1}^{i}))}{r_{\lambda}(f_{\lambda}(\xi_{t}^{I_{t+1}^{i}},y_{t+1},\varepsilon_{t+1}^{i})\mid\xi_{t}^{I_{t+1}^{i}},y_{n+1})}$
@@
(6) $\mathcal{G}_{T}(\lambda,\theta)\vcentcolon=\mathbb{E}_{Q^{N}_{\lambda,\theta}}\left[\nabla_{(\lambda,\theta)}\log\left(\prod_{t=0}^{T}\frac{1}{N}\Omega_{t}(\lambda,\theta)\right)\right],$
where $Q^{N}_{\lambda,\theta}$ now corresponds to the law of the random variables $(\varepsilon_{0}^{i})_{i=1}^{N}$ and $(\varepsilon_{t}^{i},I_{t}^{i})_{i=1}^{N}$, $t\in\{1,\ldots,T\}$, generated by the particle filter using the reparameterization trick. The approximate gradient $\mathcal{G}_{T}(\lambda,\theta)$ is indeed different from $\nabla_{(\lambda,\theta)}\mathcal{L}^{\texttt{SMC}}(\lambda,\theta)$ in that the latter contains one additional term corresponding to the expectation of the product of $\nabla_{(\lambda,\theta)}\log Q^{N}_{\lambda,\theta}$ and the logarithm of the unbiased likelihood estimator (see Le et al.,, 2018, Section A, for details). Nonetheless, since, as demonstrated by Le et al., (2018), Maddison et al., (2017) and Naesseth et al., (2018), this term generally does not make a significant contribution to the gradient and is also very difficult to estimate with reasonable accuracy, we proceed as in the mentioned works and simply discard the same.
@@ 3. Online implementation of VSMC
In this section our primary goal is to learn, for a given stream $(y_{t})_{t\geq 0}$ of observations, the proposal and model parameters as the particles evolve and new observations become available. By rewriting (6) as
(7) $\mathcal{G}_{T}(\lambda,\theta)=\mathbb{E}_{Q^{N}_{\lambda,\theta}}\left[\sum_{t=0}^{T}\nabla_{(\lambda,\theta)}\log\Omega_{t}(\lambda,\theta)\right]$
@@
(8) $\lambda_{t+1}$ $\leftarrow\lambda_{t}+\gamma_{t+1}^{\lambda}\nabla_{\lambda}\log\Omega_{t+1}(\lambda_{t},\theta_{t}),$
(9) $\theta_{t+1}$ $\leftarrow\theta_{t}+\gamma_{t+1}^{\theta}\nabla_{\theta}\log\Omega_{t+1}(\lambda_{t+1},\theta_{t}),$

### v3

@@ 1. Introduction
Being the most classical structured probabilistic generative model for serial data, state-space models (SSM), also known as general state-space hidden Markov models, are fundamental and ubiquitous in AI and statistical machine learning (Cappé et al.,, 2005; Bishop,, 2016). In SSM, any form of parameter learning or state inference typically involves the computation of complex joint posterior distributions of latent-state variables—the so-called joint-smoothing distributions—given records of observations, which is a delicate problem whose analytical solution is intractable outside the limited cases of linear Gaussian models or models with finite state space. In a recent line of research (Le et al.,, 2018; Maddison et al.,, 2017; Naesseth et al.,, 2018), this problem is addressed by combining variational inference (Blei et al.,, 2017; Kingma and Welling,, 2014) and sequential Monte Carlo (SMC) methods (Gordon et al.,, 1993; Doucet et al.,, 2001; Chopin and Papaspiliopoulos,, 2020) in order to design flexible families of variational joint-state distributions in the form of particle-trajectory laws. By optimizing the Kullback–Leibler divergence from (KLD) the law of the particle trajectories to the joint-smoothing distribution, this variational SMC (VSMC) approach is killing two birds with one stone by learning not only the unknown model parameters but also, in parallel, an optimal particle proposal kernel, the latter being a problem that has received a lot of attention in the literature (Doucet et al.,, 2000; Cornebise et al.,, 2008; Gu et al.,, 2015). The procedure can be viewed as a non-trivial extension of the importance-weighted auto-encoder (IWAE) proposed by Burda et al., (2016), where standard self-normalized importance sampling has been replaced by sequential importance sampling with systematic resampling in order to obtain a tighter evidence lower bound (ELBO). The objective of VSMC is a similar ELBO whose gradient is approximated by the expectation, under the law of the random numbers generated by the SMC algorithm, of the gradient of the logarithm of the standard particle-based likelihood estimator, the latter being unbiased (Del Moral,, 2004; Chopin and Papaspiliopoulos,, 2020), is obtained as a by-product of the particle approximation of the marginal-state posterior—or filter distribution—flow; thus, whereas traditional particle-based inference in SSM typically relies on particle approximation of the joint-smoothing distributions (with aim of approximating, e.g., the intermediate quantity of the expectation–maximization, EM, algorithm or the score function directly via the Fisher identity), which is cumbersome due to the so-called particle-path degeneracy phenomenon (Kitagawa,, 1996; Cappé et al.,, 2005, Section 8.3), VSMC allows, using the reparameterization trick, the model parameters as well as an optimal particle proposal kernel to be simultaneously learned by processing repeatedly the given data batch using a standard particle filter.
@@
In its basic form, VSMC is an offline inference technique, in the sense that it requires the full data batch to be processed between every update of the model and variational parameters. Still, in a wide range of AI and machine-learning contexts, observed data become available sequentially through a data stream, requiring learning to be performed in an online fashion. The increasing interest in online machine-learning technology also stems from the need of processing data batches of ever-increasing sizes. Thus, in this work we propose an online, stochastic approximation-based version of VSMC, online variational SMC (OVSMC), which can be used for simultaneous online model learning and proposal adaptation. On the contrary to traditional approaches to particle-based online parameter learning in SSM such as particle-based recursive maximum-likelihood (RML, Le Gland and Mevel,, 1997; Del Moral et al.,, 2015) or online EM (Mongillo and Denève,, 2008; Cappé,, 2011), which rely on particle-based online approximation of the tangent filter (filter derivative) and the EM-intermediate quantity, respectively, our approach does not involve any particle smoothing, which, in order to avoid the problem of collapsing particle ancestral genealogies, typically calls for backward-sampling techniques which can be computationally very costly. In a recent work, Campbell et al., (2021) used variational approximations of the backward kernels for online state and parameter learning. Although this, interestingly, provides a non-particle-based methodology, it is very computationally intensive (see Section 5). In addition, our method allows for effective online adaptation of the particle proposal kernel in a way that differs from typical approaches in the literature; indeed, whereas traditional approaches aim to optimize the proposal time step by time step on the basis of local criteria (see, e.g., Cornebise et al.,, 2008, 2014; Zhao et al.,, 2022), our approach is based on a global KLD-based criterion (described in detail in Section 3) which allows the particle cloud to be effectively guided toward state-space regions of high local likelihood, without running the risk of over-adapting locally the proposal to current (or temporarily neighboring) observations.
@@
Although our proposed method has similarities with the streaming variational Monte Carlo methodology proposed by Zhao et al., (2022), it is essentially different from the same. An important difference is that the aforementioned work focuses on local optimization of model and variational parameters, such that these are assumed to vary with time and are therefore optimized time step by time step, while OVSMC estimate global, amortized parameters using stochastic approximation. In addition, in order to show that our approach is statistically well founded, we provide rigorous theoretical results describing its convergence (Theorem 4.7). Under certain strong mixing assumptions, the time-normalized gradient guiding the learning of batch VSMC can, using Birkhoff’s ergodic theorem, be shown to converge as the observation batch size increases towards infinity to a deterministic function depending on the parameter as well as the particle sample size; we show that as the number of observations tends to infinity, OVSMC is solving the same problem as this ideal, ‘asymptotic’ VSMC, in the sense that the mean field targeted by OVSMC coincides with the time-normalized asymptotic gradient of VSMC.
@@
Finally, we illustrate OVSMC numerically on a number of classical SSM and more complex generative models, for which the method exhibits fast parameter learning and efficient adaptation of the particle proposal kernel. In the same numerical study, we also show that OVSMC is a strong challenger of VSMC on batch problems.
@@ 2. Background
$p_{\theta}(x_{0:t},y_{0:t})=m_{0}(x_{0})g_{\theta}(y_{0}\mid x_{0})\prod_{s=0}^{t-1}m_{\theta}(x_{s+1}\mid x_{s})g_{\theta}(y_{s+1}\mid x_{s+1}).$ (1)
@@ 2.1. Sequential Monte Carlo Methods
See Del Moral, (2004) for a comprehensive treatment of the theory of SMC. The SMC procedure consists of two core operations performed alternately: a selection step, which resamples the particles with replacement according to their weights, and a mutation step, which propagates randomly the selected particles to new locations. After importance sampling-based initialization of $(\xi_{0}^{i},\omega_{0}^{i})_{i=1}^{N}$, the random sample is updated according to Algorithm 1, in which mutation (Line 4) is executed using some generic proposal Markov transition kernel, possibly depending on $y_{t+1}$, with density $r_{\lambda}(x_{t+1}\mid x_{t},y_{t+1})$ parameterized by some vector $\lambda$ belonging to some parameter space $\mathsf{\Lambda}$. Line 3 corresponds to the selection step, where indices $(I_{t+1}^{i})_{i=1}^{N}$ guiding the resampling are drawn from the categorical distribution on $\{1,\dots,N\}$ induced by the particle weights.
@@
Determining a good proposal distribution $r_{\lambda}$ is crucial for the performance of the particle filter (see, e.g., Cornebise et al.,, 2008, 2014; Gu et al.,, 2015; Zhao et al.,, 2022), and the particles should be guided towards state-space regions of non-vanishing likelihood in order to avoid computational waste. For instance, if the latent process is diffuse while the observations are highly informative, letting naively, as in the so-called bootstrap particle filter (Gordon et al.,, 1993), $r_{\lambda}\equiv m_{\theta}$ may result in many particles being assigned a negligible weight and thus significant sample depletion. A more appealing, data-driven option is to use the locally optimal proposal satisfying $r_{\lambda}(x_{t+1}\mid x_{t},y_{t+1})\propto m_{\theta}(x_{t+1}\mid x_{t})g_{\theta}(y_{t+1}\mid x_{t+1})$ (and minimizing, e.g., the Kullback–Leibler divergence Cornebise et al., (2008)); however, this proposal is available in a closed form only in a few cases (Doucet et al.,, 2000; Cappé et al.,, 2005, Section 7.2.2.2). In Section 3 we will present a technique that allows to learn simultaneously, in an online fashion, both unknown model parameters and an efficient proposal. The proposed method is based on variational inference, which is briefly reviewed in the next section.
@@ 2.2. Variational Inference
$\mathcal{L}^{\texttt{VI}}(\lambda,\theta)\vcentcolon=\mathbb{E}_{q_{\lambda}}\left[\log\left(\frac{p_{\theta}(X_{0:T},y_{0:T})}{q_{\lambda}(X_{0:T}\mid y_{0:T})}\right)\right]\leq\log\mathbb{E}_{q_{\lambda}}\left[\frac{p_{\theta}(X_{0:T},y_{0:T})}{q_{\lambda}(X_{0:T}\mid y_{0:T})}\right]=\log p_{\theta}(y_{0:T})$ (3)
@@
The importance weighted autoencoder (IWAE, Burda et al.,, 2016) extends this idea further by optimizing a similar but improved ELBO, where the expectation is taken with respect to the law of $N$ independent and $q_{\lambda}$-distributed random variables, denoted by $q_{\lambda}^{\varotimes N}$:
$\mathcal{L}^{\tiny{\texttt{IWAE}}}(\lambda,\theta)\vcentcolon=\mathbb{E}_{q_{\lambda}^{\varotimes N}}\left[\log\left(\frac{1}{N}\sum_{i=1}^{N}\frac{p_{\theta}(X_{0:T}^{i},y_{0:T})}{q_{\lambda}(X_{0:T}^{i}\mid y_{0:T})}\right)\right]\leq\log\mathbb{E}_{q_{\lambda}^{\varotimes N}}\left[\frac{1}{N}\sum_{i=1}^{N}\frac{p_{\theta}(X_{0:T}^{i},y_{0:T})}{q_{\lambda}(X_{0:T}^{i}\mid y_{0:T})}\right]\\ =\log\mathbb{E}_{q_{\lambda}}\left[\frac{p_{\theta}(X_{0:T},y_{0:T})}{q_{\lambda}(X_{0:T}\mid y_{0:T})}\right]=\log p_{\theta}(y_{0:T}).$ (4)
@@
$\mathcal{L}^{\texttt{SMC}}(\lambda,\theta)\vcentcolon=\mathbb{E}_{Q^{N}_{\lambda,\theta}}\left[\log\left(\prod_{t=0}^{T}\frac{1}{N}\Omega_{t}\right)\right]$ ,
@@
$\omega_{t+1}^{i}(\lambda,\theta)\vcentcolon=m_{\theta}(f_{\lambda}(\xi_{t}^{I_{t+1}^{i}},y_{t+1},\varepsilon_{t+1}^{i})\mid\xi_{t}^{I_{t+1}^{i}})\frac{g_{\theta}(y_{t+1}\mid f_{\lambda}(\xi_{t}^{I_{t+1}^{i}},y_{t+1},\varepsilon_{t+1}^{i}))}{r_{\lambda}(f_{\lambda}(\xi_{t}^{I_{t+1}^{i}},y_{t+1},\varepsilon_{t+1}^{i})\mid\xi_{t}^{I_{t+1}^{i}},y_{n+1})}$ (5)
@@
$\mathcal{G}_{T}(\lambda,\theta)\vcentcolon=\mathbb{E}_{Q^{N}_{\lambda,\theta}}\left[\nabla_{(\lambda,\theta)}\log\left(\prod_{t=0}^{T}\frac{1}{N}\Omega_{t}(\lambda,\theta)\right)\right],$ (6)
where $Q^{N}_{\lambda,\theta}$ now corresponds to the law of the random variables $(\varepsilon_{0}^{i})_{i=1}^{N}$ and $(\varepsilon_{t}^{i},I_{t}^{i})_{i=1}^{N}$, $t\in\{1,\ldots,T\}$, generated by the particle filter using the reparameterization trick. The approximate gradient $\mathcal{G}_{T}(\lambda,\theta)$ is indeed different from $\nabla_{(\lambda,\theta)}\mathcal{L}^{\texttt{SMC}}(\lambda,\theta)$ in that the latter contains one additional term corresponding to the expectation of the product of $\nabla_{(\lambda,\theta)}\log Q^{N}_{\lambda,\theta}$ and the logarithm of the unbiased likelihood estimator (see Le et al.,, 2018, Appendix A, for details). Nonetheless, since, as demonstrated by Le et al., (2018), Maddison et al., (2017) and Naesseth et al., (2018), this term generally does not make a significant contribution to the gradient and is also very difficult to estimate with reasonable accuracy, we proceed as in the mentioned works and simply discard the same. Roeder et al., (2017), Tucker et al., (2018) and Finke and Thiery, (2019) discussed biased gradient approximation in the framework of the IWAE, but we are not aware of any in-depth analyses in the context of VSMC.
@@ 3. Online Implementation of VSMC
In this section our primary goal is to learn, for a given stream $(y_{t})_{t\geq 0}$ of observations, the amortized proposal and model parameters as the particles evolve and new observations become available. By rewriting (6) as
$\mathcal{G}_{T}(\lambda,\theta)=\mathbb{E}_{Q^{N}_{\lambda,\theta}}\left[\sum_{t=0}^{T}\nabla_{(\lambda,\theta)}\log\Omega_{t}(\lambda,\theta)\right]$
@@
$\lambda_{t+1}$ $\leftarrow\lambda_{t}+\gamma_{t+1}^{\lambda}\nabla_{\lambda}\log\Omega_{t+1}(\lambda_{t},\theta_{t}),$ (7)
$\theta_{t+1}$ $\leftarrow\theta_{t}+\gamma_{t+1}^{\theta}\nabla_{\theta}\log\Omega_{t+1}(\lambda_{t+1},\theta_{t}),$ (8)